tions the Beta What is the value of Risk Premium for Asset 2 What is the value of Market Risk Premium for A 3. What is the value of Beta for Asset A? 4. If Beta for Asset A decreased to 1.5, what would
Answer #1

Answer:

Asset A return Ra=10%

Market Return Rm=8%

Risk Free Return Rf=6%

1)

Risk premium for asset A= Ra-Rf=10%-6%=4%

2)

Market risk premium for Asset A=Rm-Rf=8%-6%=2%

3)

Beta = (Ra-Rf)/(Rm-Rf)=(10%-6%)/(8%-6%)=2

4)

if Beta =1.5

Then Required rate of return = Rf+Beta*(Rm-Rf)=6%+1.5*(8%-6%)=9%